The relative molecular mass of carbondioxide CO2 is 44. In an experiment, the value of γ for CO2 was found to be 1.3. Assuming that CO2 is a perfect gas, calculate the specific gas constant R and the specific heats at constant pressure cp and at constant volume, cv


1
Expert's answer
2021-07-08T09:57:46-0400

Solution.

"M=44\sdot10^{-3}kg/mol;"

"\gamma=1.3" ;

"R=\dfrac{R_0}{M};"

"R=\dfrac{8.31}{44\sdot10^{-3}}=188.9J/kgK;"

"C_V=\dfrac{R}{\gamma-1};"

"C_V=\dfrac{188.9}{1.3-1}=629.7J/kgK;"

"C_P-C_V=R\implies C_p=R+C_V;"

"C_P=188.9+629.7=818.6J\/kgK;";

Answer: "R=188.9J\/kgK;"

"C_V=629.7J\/kgK;"

"C_P=818.6J\/kgK."